OpenStudy (anonymous):

If mÐA = 3x– 2 and mÐB = 8x + 4, and the angles are complimentary, then the value of x is 9. Select one: True False

OpenStudy (anonymous):

im thinking its false ...

OpenStudy (texaschic101):

complementary angles will add up to 90 degrees. 3x - 2 + 8x + 4 = 90 -- combine like terms 11x + 2 = 90 11x = 90 - 2 11x = 88 x = 88/11 x = 8 your thinking correct.....x = 8 and not 9

OpenStudy (anonymous):

thanks (: i appreciate it .

OpenStudy (texaschic101):

no problem :)